# Learning a Neural Solver for Parametric PDE to Enhance Physics-Informed Methods

### v2

@@ 4 Experiments
We present the datasets used in the experiments in section 4.1, a comparison with selected baselines in section 4.2 and a test-time comparison with different optimizers demonstrating the remarkable effectiveness of the proposed method in section 4.3. More experiments can be found in the appendices: ablations are in appendix E and additional results and visualization in appendix F.
@@ 4.1 Datasets
We consider several representative parametric equations for our evaluation. More details about the data generation are presented in appendix D. Our objective is to learn a neural solver able to solve accurately and fast a new instance of a PDE, given its parametric form, and the values of the parameters $\gamma$, forcing terms $f$ and initial/boundary conditions $g$, i.e. $(\gamma,f,g)\mapsto u$. Solving is performed with a few iterations of the neural solver (algorithm 1). For that, one trains the neural solver on a sample of the PDE parameter instances, see table 1 for the parameter distributions used for each parametric PDE. Evaluation is performed on unseen sets of parameters within the same PDE family.
@@
Helmholtz: We generate a dataset following the $1d$ static Helmholtz equation $u^{\prime\prime}(x)+\omega^{2}u(x)=0$ with boundary conditions $u(0)=u_{0}\text{ and }u^{\prime}(0)=v_{0}$. We generate $1,024$ trajectories with varying $\omega$, $u_{0}$ and $v_{0}$ with a spatial resolution of $256$. Poisson: We generate a dataset following the $1d$ static Poisson equation with forcing term: $-u^{\prime\prime}(x)=f(x)$ with $u(0)=u_{0}$ and $u^{\prime}(0)=v_{0}$. The forcing term $f$ is a periodic function, $f(x)=\frac{\pi}{K}\sum_{i=1}^{K}a_{i}i^{2r}\sin(\pi x)$, with $K=16$ and $r=-0.5$. We generate $1,000$ trajectories with varying $u_{0},v_{0}$ and $f$ (through changing $a_{i}$) with a spatial resolution of $64$. Reaction-Diffusion: In Krishnapriyan et al. (2021); Toloubidokhti et al. (2024), the authors propose a non-linear reaction-diffusion (NLRD). This PDE has been shown to be a failure case for PINNs (Krishnapriyan et al., 2021). We generate $1,000$ trajectories by varying the parameters of the PDE: $\nu$ and $\rho$ (see table 1). Spatial resolution is $256$ and temporal resolution is $100$. The PDE is solved on $[0,1]^{2}$. Darcy Flow: The $2d$ Darcy Flow dataset is taken from (Li et al., 2020) and commonly used in the operator learning literature (Li et al., 2023b; Goswami et al., 2022). For this dataset, the forcing term $f$ is kept constant $f=1$, and $a(x)$ is a piece-wise constant diffusion coefficient taken from (Li et al., 2020). We kept $1,000$ trajectories (on the $5,000$ available) with a spatial resolution is $64\times 64$. Heat: The $2d+t$ Heat equation is simulated as proposed in (Zhou & Farimani, 2024). For this dataset, the parameter $\nu$ is sampled from $\mathcal{U}[2\times 10^{-3},2\times 10^{-2}]$ and initial conditions are a combination of sine functions with varying number of terms, amplitude and phase. A summary of the datasets and the variable parameters for each dataset are presented in table 1 and more details on the dataset are provided in appendix D.
Experiments have been conducted on NVIDIA TITAN V ($12$ Go) for $1d$ datasets and NVIDIA RTX A6000 GPU with 49Go for $1d$ + time or $2d$ datasets. For all datasets, $800$ PDEs are considered during training and $200$ for testing. All metrics reported are evaluated on test samples (i.e. PDEs not seen during training. Coefficient as well as initial and/or boundary condition can vary from training).
@@ 4.2 Comparison with baselines
We performed comparisons with several baselines including fully data-driven supervised approaches trained from a data-loss only, unsupervised methods relying only on a PDE loss, and hybrid techniques trained from PDE + DATA loss. Networks size and training details are described in appendix C. In this experiment, we considered training the models using the training sets (physical losses or MSE when possible) unless stated otherwise.
@@
Fully-supervised We train a standard MLP to learn the mapping $(\gamma,f,g)\mapsto\Theta$, using as loss function $\mathcal{L}_{\text{DATA}}=\mathbb{E}_{\gamma,f,g,u}\left[||u_{\Theta_{L}}-u_{\gamma,f,g,u}||\right]$ with $u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)$, the $\psi_{i}(.)$ being fixed B-Spline basis functions (see appendix C). We denote this baseline as MLP+basis. Unsupervised We compare our approach with unsupervised physics-informed models (Raissi et al., 2019). While the initial version of PINNs solves only one PDE instance at a time and requires retraining for each new instance, we developed here a parametric version of PINNs (PPINNs) where the PDE parameters are fed to the network. A similar approach has been used by (Zhang et al., 2023). In addition to PINNs, we also compare our method to the Physic-informed DeepONet (designed as PO-DeepONet for Physics-Only DeepONet) (Wang et al., 2021), which is designed to learn an operator for function-to-function mappings from physical losses and handles parametric PDEs. The mapping learned for the two unsupervised baselines is $(x,\gamma,f,g)\mapsto u(x)$. In order to provide a fair comparison with our optimization method, we fine-tuned the unsupervised baselines for each specific PDE instance for a few steps ($10$ or $20$). Comparison to pre-conditionning We compare our approach with vanilla PINNs (Raissi et al., 2019), i.e. by fitting one PINN per PDE in the test set and averaging. We optimize the PDE losses using L-BFGS (Liu & Nocedal, 1989) and refer to this baselines as PINNs+L-BFGS. As discussed in (Rathore et al., 2024), L-BFGS can be considered as a non linear preconditioning method for the linear Physics-Informed methods and fastens convergence. For this baseline one model is trained and evaluated for each PDE in the test set. We report the reader to appendix C for more details on the training procedure. Hybrid Finally, we compare with neural operators, i.e., trained to learn mappings $(x,\gamma,f,g)\mapsto u(x)$ using a combination of physical and data loss: $\mathcal{L}_{\textnormal{DATA}}+\mathcal{L}_{\textnormal{PDE}}$. We use as baselines Physics-Informed Neural Operator (PINO) (Li et al., 2023b) and Physics-Informed DeepONet (PI-DeepONet) (Goswami et al., 2022). As already indicated, for a fair comparison, the Unsupervised and Hybrid baselines, are fine-tuned on each specific PDE instance for a few steps ($10$ on all dataset except for Heat for which $20$ steps are made). Ours We represent the solution $u_{\Theta}$ with a linear combination of B-Spline functions for $\Psi$ (Piegl & Tiller, 1996). This was motivated by the nature of B-Splines which allows to capture local phenomena. However, other bases could be used such as Fourier, Wavelet or Chebychev Polynomials. More details about the construction of the B-Spline basis are available in appendix C. The neural solver $\mathcal{F}_{\varrho}$ is composed of Fourier Layers (Li et al., 2020) that allow us capturing the range of frequencies present in the phenomenon.
The proposed method and all the baselines are implemented with PyTorch (Ansel et al., 2024).
@@
Unsupervised PPINNs 9.89e-1 4.30e-2 3.94e-1 8.47e-1 1.27e-1
PINNs+L-BFGS 9.86e-1 8.83e-1 6.13e-1 9.99e-1 9.56e-1
@@
Table 2: Results of trained models - metrics in Relative MSE on test set. Best performances are highlighted in bold, second best are underlined.
Results:
Table 2 presents the comparison with the baselines. Each method is trained to solve a parametric equation, using a sample of the equation parameters. The evaluation set is composed of several PDE instances sampled from unseen PDE parameters $(\gamma,f,g)$. The proposed method is ranked first or second on all the evaluations. The most comparable baselines are the unsupervised methods, since at inference they leverage only the PDE residual loss, as our method does. Therefore our method should be primarily compared toe these baselines. Supervised and hybrid methods both incorporate a data loss and make different assumptions while solving a different optimization problem.
Table 2 clearly illustrates that unsupervised Physics-informed baselines all suffer from ill-conditioning and do not capture the dynamics. Compared to these baselines, the proposed method improves at least by one order of magnitude on all cases. This is observed for PPINNs as well as on PINNs fitted on one equation only. PINNs baseline performs poorly on these datasets because of the ill-conditioning nature of the PDE, requiring numerous optimization steps to achieve accurate solutions (appendix C). We observe that our neural solver has better convergence properties than other Physics-Informed methods. At will be seen later it also converges much faster.
The supervised baseline performs well on all the PDEs except Poisson and Heat. The data loss used for training this model is the mean square error which is well-behaved and does not suffer from optimization problems as the PDE loss does. We note that our method reaches similar or better performances on every dataset, while relying only on physical information at inference (algorithm 1) and solving a more complex optimization problem.
The hybrid approaches, do not perform well despite taking benefits from the PDE+DATA loss and from adaptation steps at test time. Again, the proposed method is often one order of magnitude better than the hybrids except on NLRD, where it has comparable performances. This shows that the combination of physics and data losses is also hard to optimize suffering from ill-conditioning.
@@ 4.3 Optimization for solving new equations
The main motivation for our learned PDE solver is to accelerate the convergence to a solution, w.r.t. predefined solvers, for a new equation. In order to assess this property, we compare the convergence speed at test time inference with classical solvers, PINNs and pre-trained PINO as detailed below. Results are presented in fig. 2 for the Poisson equation with performance averaged on $20$ new instances of the Poisson equation. The experiments for the other datasets are provided in appendix F.
@@
Baseline optimizers As for the classical optimizer baselines, we used SGD, Adam (Kingma & Ba, 2015) and L-BFGS (Krishnapriyan et al., 2021). These optimizers are used to learn the coefficient of the B-Spline basis expansion in the model $u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)$. This provides a direct comparison to our iterative neural solver. PINNs - We also compare to the standard PINNs (Raissi et al., 2019), i.e. by fitting one Neural Network (NN) per equation. Note that this requires full training from scratch for each new equation instance and this is considerably more computationally demanding than solving directly the parametric setting. This baseline is similar to the Adam optimizer mentioned above, except that the ansatz for this experience is a multilayer perceptron instead of a linear combination of B-Spline basis. Hybrid pre-training strategies - Finally we compare against the hybrid PINO pre-trained on a set of several parametric PDE instances and then fine-tuned on a new PDE instance using only the PDE loss associated to this instance. Ours - We train our model as explained in algorithm 2 and show here the optimization process at test-time. In order to perform its optimization, our model leverages the gradient of the physical loss and the PDE parameters (coefficients, initial/boundary conditions). In this experiment, we use $L=5$ steps for a better visualization (whereas, we used $L=2$ in table 2).
Result: fig. 2 compares the number of optimization steps required for the different methods. Our neural solver converges very fast in only a few steps (5 here) to a good approximation of the solution, while all the other methods require thousands of iterations - we stopped here at $10,000$ steps. The classical optimizers (SGD, Adam, L-BFGS) do not converge for a new equation. The baseline PINNs trained here from scratch on each new equation shows an erratic convergence behavior. Pretrained PINO behaves better than the other baselines but still did not converged after $10,000$ steps. This clearly demonstrates the potential of our learned solver to deal with physical losses w.r.t. alternative pre-defined solvers.
@@ 5 Conclusion
We have presented a learning based PDE solver that allows fast test-time optimization of physical losses. Our method succeeds to considerably accelerate the optimization process for the complex problem of minimizing Physical losses, and is several orders of magnitude faster than classical hand-defined optimization methods such as Adam or L-BFGS.
Limitations and Future Work
The proposed method although efficient, could still be improved in several ways leading to further exploration of this idea. First, the memory needed to train our iterative algorithms can be more demanding than for usual ML model learning. Indeed, back-propagating through the iterations adds complexity and this might become an issue on higher dimensional basis. Second, we focused here on solution approximations expressed as a linear expansion in a pre-defined bases. More expressive representations such as neural networks could be investigated, although our preliminary experiments have shown even more ill-conditioning, due to the compositional nature of the neural network. More sophisticated training schemes, optimizers, or initializations could enhance the performances of the optimization process. Future work will explore such issues to improve its scalability and capabilities for broader application of the proposed method.
@@
Solving PDE is of crucial interest in many applications of science and engineering. While we do not directly target such real-world applications in this paper, one should acknowledge that solvers can be used in various ranges of scenarios including weather, climate, medical, aerodynamics, and military applications.
Reproducibility Statement
We will provide all elements that can be necessary to reproduce the experiments. Hyper-parameters, baselines configurations and training details are shown in appendices C, 4, 3, 2 and 1. The generation of the datasets are explained in appendices D, 1 and 5 with the solver used or the source, as well as the parameter ranges considered in the experiments. Finally, we provide a theoretical analysis of the model, under ideal scenario in appendix B. Code will be released upon acceptance as well as scripts to generate data when needed.

### v3

@@ 4 Experiments
We present the datasets used in the experiments in section 4.1, a comparison with selected baselines in section 4.2, and a test-time comparison with different optimizers demonstrating the remarkable effectiveness of the proposed method in section 4.3. Finally, we make a comparison of the training and inference time in section 4.4. Experimental details and additional experiments can be found in the appendices: ablations are in appendix E and additional results and visualization are in section E.6, section E.7, and appendix F.
@@ 4.1 Datasets
We consider several representative parametric equations for our evaluation. More details about the data generation are presented in appendix C. Our objective is to learn a neural solver able to solve quickly and accurately a new instance of a PDE, given its parametric form, and the values of the parameters $\gamma$, forcing terms $f$ and initial/boundary conditions $g$, i.e. $(\gamma,f,g)\mapsto u$. Solving is performed with a few iterations of the neural solver (algorithm 1). For that, one trains the neural solver on a sample of the PDE parameter instances, see table 1 for the parameter distributions used for each parametric PDE. Evaluation is performed on unseen sets of parameters within the same PDE family. Helmholtz: We generate a dataset following the $1d$ static Helmholtz equation $u^{\prime\prime}(x)+\omega^{2}u(x)=0$ with boundary conditions $u(0)=u_{0}\text{ and }u^{\prime}(0)=v_{0}$. We generate $1,024$ trajectories with varying $\omega$, $u_{0}$, and $v_{0}$ with a spatial resolution of $256$. Poisson: We generate a dataset following the $1d$ static Poisson equation with forcing term: $-u^{\prime\prime}(x)=f(x)$ with $u(0)=u_{0}$ and $u^{\prime}(0)=v_{0}$. The forcing term $f$ is a periodic function, $f(x)=\frac{\pi}{K}\sum_{i=1}^{K}a_{i}i^{2r}\sin(\pi x)$, with $K=16$ and $r=-0.5$. We generate $1,000$ trajectories with varying $u_{0},v_{0}$, and $f$ (through changing $a_{i}$) with a spatial resolution of $64$. Reaction-Diffusion: In Krishnapriyan et al. (2021); Toloubidokhti et al. (2024), the authors propose a non-linear reaction-diffusion (NLRD). This PDE has been shown to be a failure case for PINNs (Krishnapriyan et al., 2021). We generate $1,000$ trajectories by varying the parameters of the PDE: $\nu$ and $\rho$ (see table 1). Spatial resolution is $256$ and temporal resolution is $100$. The PDE is solved on $[0,1]^{2}$. Darcy Flow: The $2d$ Darcy Flow dataset is taken from (Li et al., 2020) and is commonly used in the operator learning literature (Li et al., 2023b; Goswami et al., 2022). For this dataset, the forcing term $f$ is kept constant $f=1$, and $a(x)$ is a piece-wise constant diffusion coefficient taken from (Li et al., 2020). We kept $1,000$ trajectories (on the $5,000$ available) with a spatial resolution is $64\times 64$. Heat: The $2d+t$ Heat equation is simulated as proposed in (Zhou & Farimani, 2024). For this dataset, the parameter $\nu$ is sampled from $\mathcal{U}[2\times 10^{-3},2\times 10^{-2}]$ and initial conditions are a combination of sine functions with a varying number of terms, amplitude, and phase. A summary of the datasets and the varying parameters for each PDE are presented in table 1 and more details on the dataset are provided in appendix C. Experiments have been conducted on NVIDIA TITAN V ($12$ Go) for $1d$ datasets to NVIDIA RTX A6000 GPU with 49Go for $1d$ + time or $2d$ datasets. For all datasets, $800$ PDEs are considered during training and $200$ for testing. All metrics reported are evaluated on test samples (i.e. PDEs not seen during training. Coefficients as well as initial and/or boundary conditions can vary from training).
@@ 4.2 Comparison with baselines
We performed comparisons with several baselines including fully data-driven supervised approaches trained from a data-loss only, unsupervised methods relying only on a PDE loss, and hybrid techniques trained from PDE + DATA losses. Network size and training details are described in appendix D. In this experiment, we considered training the models using the training sets (physical losses or MSE when possible) unless stated otherwise.
@@
Fully supervised We train a standard MLP to learn the mapping $(\gamma,f,g)\mapsto\Theta$, using as loss function $\mathcal{L}_{\text{DATA}}=\mathbb{E}_{\gamma,f,g,u}\left[||u_{\Theta_{L}}-u_{\gamma,f,g}||\right]$ with $u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)$, the $\psi_{i}(.)$ being fixed B-Spline basis functions (see appendix D). We denote this baseline as MLP+basis. Unsupervised We compare our approach with unsupervised physics-informed models (Raissi et al., 2019). While the initial version of PINNs solves only one PDE instance at a time and requires retraining for each new instance, we developed here a parametric version of PINNs (PPINNs) where the PDE parameters are fed to the network (similarly to (Zhang et al., 2023)). Finally, we used (Cho et al., 2024)’s (P2INNs) method as a physics-informed baseline specifically designed for parametric PDEs. In addition to PINNs-methods, we also compare our solver to the Physic-informed DeepONet (PO-DeepONet for Physics-Only DeepONet) (Wang et al., 2021b), which is designed to learn an operator for function-to-function mappings from physical losses and handles parametric PDEs. The mapping learned for the two unsupervised baselines is $(x,\gamma,f,g)\mapsto u_{\gamma,f,g}(x)$. In order to provide a fair comparison with our optimization method, we fine-tuned the unsupervised baselines for each specific PDE instance for a few steps ($10$ or $20$). Comparison to preconditioning We compare our approach with vanilla PINNs (Raissi et al., 2019), i.e. by fitting one PINN per PDE in the test set and averaging the final errors. We optimize the PDE losses using L-BFGS (Liu & Nocedal, 1989) and refer to this baseline as PINNs+L-BFGS. As discussed in (Rathore et al., 2024), L-BFGS can be considered as a nonlinear preconditioning method for Physics-Informed methods and fastens convergence. Finally, we use the training strategy proposed by (Rathore et al., 2024) i.e. trained PINNs using successive optimizer (Adam + L-BFGS). This baseline is denoted as PINNs-multi-opt. For these baselines, one model is trained and evaluated for each PDE in the test set. We report the reader to appendix D for more details on the training procedure. Hybrid Finally, we compare our proposed method with neural operators, i.e., models trained to learn mappings $(x,\gamma,f,g)\mapsto u_{\gamma,f,g}(x)$ using a combination of physical and data loss: $\mathcal{L}_{\textnormal{DATA}}+\mathcal{L}_{\textnormal{PDE}}$. We use as baselines Physics-Informed Neural Operator (PINO) (Li et al., 2023b) and Physics-Informed DeepONet (PI-DeepONet) (Goswami et al., 2022). As already indicated, for a fair comparison, the Unsupervised and Hybrid baselines are fine-tuned on each specific PDE instance for a few steps ($10$ on all datasets except for Heat for which $20$ steps are made). Ours We represent the solution $u_{\Theta}$ with a linear combination of B-Spline functions for $\Psi$ (Piegl & Tiller, 1996). This was motivated by the nature of B-Splines which allows to capture local phenomena. However, other bases could be used such as Fourier, Wavelet or Chebychev Polynomials. The neural solver $\mathcal{F}_{\varrho}$ is composed of Fourier Layers (FNO) (Li et al., 2020) that allow us to capture the range of frequencies present in the phenomenon. We refer the reader to appendix D for more details about the construction of the B-Spline basis and the training hyper-parameters.
@@
Unsupervised PINNs+L-BFGS 9.86e-1 8.83e-1 6.13e-1 9.99e-1 9.56e-1
PINNS-multi-opt 8.47e-1 1.18e-1 7.57e-1 8.38e-1 6.10e-1
PPINNs 9.89e-1 4.30e-2 3.94e-1 8.47e-1 1.27e-1
P2INNs 9.90e-1 1.50e-1 5.69e-1 8,38e-1 1.78e-1
@@
Table 2: Results of trained models - metrics in Relative MSE on the test set. Best performances are highlighted in bold, and second best are underlined.
Results: Table 2 presents the comparison with the baselines. We recall that the evaluation set is composed of several PDE instances sampled from unseen PDE parameters $(\gamma,f,g)$. The proposed method is ranked first or second on all the evaluations. The most comparable baselines are the unsupervised methods, since at inference they leverage only the PDE residual loss, as our method does. Therefore our method should be primarily compared to these baselines. Supervised and hybrid methods both incorporate data loss and make different assumptions while solving a different optimization problem.
Table 2 clearly illustrates that unsupervised Physics-informed baselines all suffer from ill-conditioning and do not capture the dynamics. Compared to these baselines, the proposed method improves at least by one order of magnitude in all cases. PINNs baseline performs poorly on these datasets because of the ill-conditioning nature of the PDE, requiring numerous optimization steps to achieve accurate solutions (appendix D). This is observed on PINNs models for parametric PDEs (PPINNs and P2INNs) as well as on PINNs fitted on one equation only (PINNs+L-BFGS and PINNs-multi-opt). We observe that our neural solver has better convergence properties than other Physics-Informed methods. As will be seen later it also converges much faster.
The supervised baseline performs well on all the PDEs except Poisson and Heat. The data loss used for training this model is the mean square error which is well-behaved and does not suffer from optimization problems as the PDE loss does. We note that our method reaches similar or better performances on every datasets, while relying only on physical information at inference (algorithm 1) and solving a more complex optimization problem.
The hybrid approaches, do not perform well despite taking benefits from the PDE+DATA loss and from adaptation steps at test time. Again, the proposed method is often one order of magnitude better than the hybrids except on NLRD, where it has comparable performances. This shows that the combination of physics and data losses is also hard to optimize, and suffers from ill-conditioning.
@@ 4.3 Optimization for solving new equations
The main motivation for our learned PDE solver is to accelerate the convergence to a solution, w.r.t. predefined solvers, for a new equation. In order to assess this property, we compare the convergence speed at test time inference with classical solvers, PINNs, and pre-trained PINO as detailed below. Results are presented in fig. 2 for the Poisson equation with performance averaged on $20$ new instances of the Poisson equation. This experiment is also performed on the other datasets in appendix F.
Baseline optimizers As for the classical optimizer baselines, we used SGD, Adam (Kingma & Ba, 2015), and L-BFGS (Krishnapriyan et al., 2021). These optimizers are used to learn the coefficient of the B-Spline basis expansion in the model $u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)$. This provides a direct comparison to our iterative neural solver. PINNs - We also compare to the standard PINNs (Raissi et al., 2019), i.e. by fitting one Neural Network (NN) per equation. Note that this requires full training from scratch for each new equation instance and this is considerably more computationally demanding than solving directly the parametric setting. This baseline is similar to the Adam optimizer mentioned above, except that the ansatz for this experience is a multilayer perceptron instead of a linear combination of a B-Spline basis. Hybrid pre-training strategies - Finally we compare against the hybrid PINO pre-trained on a set of several parametric PDE instances and then fine-tuned on a new PDE instance using only the PDE loss associated to this instance. Ours - We train our model as explained in algorithm 2 and show here the optimization process at test time. In order to perform its optimization, our model leverages the gradient of the physical loss and the PDE parameters (coefficients, initial/boundary conditions). In this experiment, we use $L=5$ steps for a better visualization (whereas, we used $L=2$ in table 2).
@@
Result: fig. 2 compares the number of optimization steps required for the different methods. Our neural solver converges very fast in only a few steps (5 here) to a good approximation of the solution, while all the other methods require thousands of iterations - we stopped here at $10,000$ steps. The classical optimizers (SGD, Adam, L-BFGS) do not converge for a new equation. The baseline PINNs trained here from scratch on each new equation show an erratic convergence behavior. Pretrained PINO behaves better than the other baselines but still did not converged after $10,000$ steps. This clearly demonstrates the potential of our learned solver to deal with physical losses w.r.t. alternative pre-defined solvers.
4.4 Computational time
An important aspect of solving PDEs is the computational time required for each solution. Methods should find a trade-off between achieving high performance and maintaining reasonable computational costs. In section E.3, we provide the training (table 16) and inference (table 17) times for our method compared to various baselines. Our results show that while our method has comparable training and inference times to other approaches, it demonstrates substantially better precision (table 2).
@@ 5 Conclusion
We have presented a PDE solver learned from data that allows fast test-time optimization of physical losses. Our method succeeds to considerably accelerate the optimization process for the complex problem of minimizing physical losses and is several orders of magnitude faster than classical hand-defined optimization methods such as Adam or L-BFGS. Limitations and Future Work While efficient, the proposed method can be further improved. First, training our iterative algorithms requires more memory than standard machine learning models due to the complexity of backpropagation through iterations, which becomes challenging in higher-dimensional bases. Second, we have focused on solution approximations expressed as a linear expansion in predefined bases. More expressive representations, such as neural networks, could be explored; however, our preliminary experiments indicate increased ill-conditioning due to the compositional nature of neural networks. More sophisticated training schemes could enhance the optimization process. Future work will investigate these directions to improve scalability and broaden the applicability of the proposed method. Reproducibility Statement Hyper-parameters, baselines configurations and training details are detailed in appendices D, 5, 4, 2 and 1. The creation of the datasets is explained in appendices C, 1 and 3. Finally, we provide a theoretical analysis of the model, under an ideal scenario in appendix B. Code is available at https://github.com/2ailesB/neural-parametric-solver.
@@
Solving PDE is of crucial interest in many applications of science and engineering. While we do not directly target such real-world applications in this paper, one should acknowledge that solvers can be used in various ranges of scenarios including weather, climate, medical, aerodynamics, industry, and military applications.